I am helping my daughter and trying to find the rule on determining the least common denominator. This is from worksheets in Algebra I but doesn't involve a book yet. I am trying to foind the LCD with two different fractions.
Thanks.
Mark Answer by stanbon(75887) 
 Example:
Let's say you have 1/24 + 1/75 and want the lcd.
---------------------
Find the prime-factor form of each denominator:
24 = 2^3*3
75 = 3*5^2
----------------
The lcd must contain each of the prime numbers to the highest power present
in the denominators.
------------------
So the lcd must have 2,3,and 5 to the highest power present in the
denominators.
----------------------------
the lcd is (2^3)* (3) * (5^2) = 8*3*25 =
---------------------------------
So, does that work?
The problem was 1/24 + 1/75
Rewrite each fraction with the lcd as its denominator:
Divide 24 into 600 to get 25.
Multiply numerator and denominator of the 1st fraction by 25 to get:
25/600
-------------
Divide 75 into 600 to get 8
Multiply numerator and denominator of the 2nd fraction by 8 to get:
8/600
----------
Add the two resulting fractions to get:
1/24 + 1/75 = 25/600 + 8/600 = 33/600 = 11/200
=====================================
Cheers,
Stan H.